Tutor,
If I could bother you for some help solving this problem I would appreciate it.
Find the x intercepts:
y = x^2 + x - 6
Thank you.

Answer by stanbon(75887)

Find the x intercepts:
y = x^2 + x - 6
------------------
Let y = 0 and solve for x:
x^2+x-6=0
Factor:
(x+3)(x-2)=0
x+3 = 0 or x-2 = 0
x = -3 or x = 2
--------------------
Cheers,
Stan H.